OpenStudy (anonymous):

the perimeter of a triangle is 47 centimeters. The longest side is 7 centimeters less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of all sides.

OpenStudy (anonymous):

Why u post it again?

OpenStudy (anonymous):

bc i dont understand still how you got the answer

OpenStudy (anonymous):

i do not understand your explanation

OpenStudy (anonymous):

Sides, x , y and x+y-7 add up to 47...that the bit u didn't understand?

OpenStudy (anonymous):

let shortest = x, longest = y, the thrid side = z y=x+z-7 (longest side is 7 cm less than sum of x and z) 2x + 4 = y (twice the shortest is 4 cm less than longest) the perimeter is 47, so x + y + z = 47 you now have system of equations 3 equations and 3 unknowns...can u take it from there?

OpenStudy (anonymous):

the equations x+y+z = 47 x+z-7 = y 2x+ 4 = y

OpenStudy (anonymous):

do you want me to solve the whole thing?

OpenStudy (anonymous):

I already did it, he didn't get it...

OpenStudy (anonymous):

2x+2y-7=47 2x+4 = x+y-7 -> x= y-11 etc... (ie sub in first one)

OpenStudy (anonymous):

im trying to understand it . i am having a hard time. i dont want someone to give me the answer i need to undetstand it for myself

OpenStudy (anonymous):

ok but do you understand how i set up the problem? with the system of equations i would like to know what you dont understand so i can help you more, are you not familiar with how to solve system of equations? or you just have hard time set the problem up?

OpenStudy (anonymous):

I got it. thanks for helping